Three identical very dense masses of 6000 kg each are placed on the x axis. One mass is at x1 = -120 cm , one is at the origin, and one is at x2 = 450 cm .

Q:

What is the magnitude of the net gravitational force Fgrav on the mass at the origin due to the other two masses?

Take the gravitational constant to be G = 6.67×10−11 N⋅m2/kg2 .

Q:

What is the direction of the net gravitational force on the mass at the origin due to the other two masses?
Answer #1

1. Gravitational force is attractive.

Fg = G M m / r^2

dueto x1:

Fg1 = - (6.67 x 10^-11 x 6000 x 6000 ) / (1.20^2) = - 1.6675 x 10^-3 N

due to x2 :

Fg2 = (6.67 x 10^-11 x 6000 x 6000 ) / (4.5^2) =   0.119 x 10^-3 N

Fnet = Fg1 + Fg2 = - 1.548 x 10^-3

magnitude = 1.548 x 10^-3 N


2.
direction = toward -ve x axis.